OpenStudy (anonymous):

Find the value of f(x)=3x^2-x+23 for x=-4

OpenStudy (haseeb96):

it is so easy if u solve it as i do here in the question it is given to u that x=-4 okay ? now the equation is f(x)=3x^2 - x +23 okay now put the x=-4 in the above equation on the place of ''x'' okay? so it would be f(-4) = 3 (-4)^2 -(-4) +23 = 3 (16) + 4+23 since (-4)^2 =16 and - (-4)= +4 = 48 + 4 +23 f(-4) =75 this is your correct answer if u have any question then u can ask to me